Write each ratio or rate in simplest form. 27 to 36 1. 1. Math 2. Math 10 yards to 75 yards 2. 234 miles on 9 gallons of gas9. Solve. 9. If you can buy 25 pounds of dog food for $7.25 or 40 pounds for $11.20, which would be the better buy? Math 10.Rewrite. 48% as a fraction in simplest form 12. 12. Math 13. Math 0.5% as a decimal 13. Check 14. Math 2.15 as a percent 14.17. Math What percent of 6 is 12 ? 17. Check Math 28 Chapter 5&6 Test, Form A - Page 18. 40% of what number is 12 ? 18. Math20a. 20. Jaynes salary last year was $52,000. If she receives a 3% raise this year, what will be the dollar amount of her ra
Answer #1


0 27 to 36 27 36 mit log arda to 75 yards Like 75 234 miles on 9 gallon of gas @ 234 239 26 miles/gallon 4 134 feet in 5 Sec-
Answer #2

First 10 questions answered :

ANSWERS :


1.  27 to 36  = 3 to 4 


2. 18 to 75 = 6 to 25 


3. 234 miles on 9 gallons of gas = 26 miles per gallon 


4. 134 ft in 5 secs = 26.8 ft per sec 


9.


25 pounds dog food for $7.25 = 1 pound dog food for $0.29 


40 pounds dog food for $11.20 = 1 pound dog food for $0.28 


So, the latter one is cheaper and should be bought. (ANSWER).


10.


Decrease cost  to original cost 

= (28000 - 21000) / 28000

= 7000 / 28000

= 1/4 (ANSWER).

answered by: Tulsiram Garg